# Damper-B-PINN: Damper Characteristics-Based Bayesian Physics-Informed Neural Network for Vehicle State Estimation

### v2

@@ Abstract
Accurate state estimation is fundamental to intelligent vehicles. Wheel load, one of the most important chassis states, serves as an essential input for advanced driver assistance systems (ADAS) and exerts a direct influence on vehicle stability and safety. However, wheel load estimation remains challenging due to the complexity of chassis modeling and the susceptibility of nonlinear systems to noise. To address these issues, this paper first introduces a refined suspension linkage-level modeling approach that constructs a nonlinear instantaneous dynamic model by explicitly considering the complex geometric structure of the suspension. Building upon this, we propose a damper characteristics-based Bayesian physics-informed neural network (Damper-B-PINN) framework to estimate dynamic wheel load, which leverages the suspension dynamics as physical guidance of PINN while employing Bayesian inference to mitigate the effects of system noise and uncertainty. Moreover, a damper-characteristic physics conditioning (DPC) module is designed for embedding physical prior. The proposed Damper-B-PINN is evaluated using both high-fidelity simulation datasets generated by CarSim software and real-world datasets collected from a Formula Student race car. Experimental results demonstrate that our Damper-B-PINN consistently outperforms existing methods across various test conditions, particularly extreme ones. These findings highlight the potential of the proposed Damper-B-PINN framework to enhance the accuracy and robustness of dynamic wheel load estimation, thereby improving the reliability and safety of ADAS applications.
@@ I Introduction
State estimation in real-world engineering systems subject to noise, such as wheel load estimation in vehicle chassis systems, remains a longstanding challenge due to the complex and imperfect nonlinear relationships between inputs and outputs [1, 2]. Traditional methods like EKF sometimes fail to realize accurate estimation in real-world scenarios [3]. In recent years, data-driven methods have increasingly been employed to approximate closures in nonlinear models and to estimate parameters and functional relationships within mathematical frameworks [4, 5]. Among these, physics-informed neural networks (PINNs) have emerged as a powerful paradigm for both forward (inference) and inverse (identification) problems due to their ability to integrate physical laws into the learning process [6, 7].
Figure 1: Testing platform (A Formula Student racing car) and standard testing circuits.
Recent studies on PINNs have explored their applications in various complex engineering fields, such as robotics [8], automotive [9], and transportation [10], demonstrating strong performance across diverse tasks. Additionally, novel optimization methods have been proposed to enhance the accuracy and generalization of PINNs [11]. Despite the fact that PINNs have demonstrated success in integrating physics into deep neural networks (DNNs) framework, enabling the concurrent use of physics as explicit knowledge and data as implicit knowledge, they lack built-in uncertainty quantification, limiting their applicability in scenarios with high noise levels.
@@
Recently, physics-informed BNNs have been introduced as a hybrid approach that integrates PINNs and Bayesian techniques to quantify uncertainties in both data and models [14]. Considering sparse and noisy measurements in boundary conditions and source terms in real-world applications, Yang et al. [15] solved both forward and inverse problems involving linear and nonlinear PDEs with noisy data. In the wheel load estimation task, there are two main challenges: (1) Inaccurate dynamic modeling will mislead the PINN, ultimately resulting in the physical model failing to match the data distribution. (2) The chassis sensors have significant noise, and traditional PINN is only constrained through the physical loss, making it fail to stably embed the physical information during the training process.
To address these problems, we design a damper characteristics-based Bayesian physics-informed neural network (Damper-B-PINN), a BNN framework based on damper properties and introduce a physical model as a priori knowledge within the network structure to enhance accuracy in system estimation, as shown in Figure 2. The main contributions of this paper are summarized as follows:
  • A Bayesian PINN framework with a damper-characteristic physics conditioning (DPC) module, which is introduced to reduce the impact of environmental noise on systems with physical prior, enhancing the robustness of networks.
II Methodology
II-A Problem Formulation
@@
II-B Suspension Linkage-Level Modeling
Since vehicle suspensions typically consist of multiple connecting rods forming complex single-degree-of-freedom (1 DOF) or two-degree-of-freedom (2 DOF) structures, their geometric parameters evolve dynamically during driving. For computational simplicity, the mass of individual suspension links is neglected, and only the unsprung mass $m_{u}$, which includes the wheel and steering knuckle, is considered. The hinge joints of the suspension system, commonly referred to as “hard points”. Detailed components are labeled in Figure 3. In this study, the hard points are defined as follows: $\mathbb{u_{1}}$ and $\mathbb{u_{2}}$ (front and rear points of the upper control arm, UCA); $\mathbb{l_{1}}$ and $\mathbb{l_{2}}$ (front and rear points of the lower control arm, LCA); $\mathbb{p_{1}}$ and $\mathbb{p_{2}}$ (lower and upper joints of the spring-damper assembly); $\mathbb{t}$ (hinge joint between the steering knuckle and steering rack); and $\mathbb{s_{1}}$, $\mathbb{s_{2}}$, and $\mathbb{s_{3}}$ (upper, front, and lower points of the steering knuckle).
@@ II-B Suspension Linkage-Level Modeling
where $f$ represents all the solving processes from Equation (LABEL:eq2) to Equation (7). Then, the complex suspension dynamics will be integrated into the BNN as physical information.
II-C Bayesian Physics-Informed Neural Network
II-C1 Physics-Informed Bayesian Framework
@@
II-C2 Variational Inference with Physics-Informed Likelihood
@@
II-C3 Normal-Sigmoid-Dropout Method
@@
II-D Damper-Characteristic Physics Conditioning Embedding
@@
The DPC encoder is designed to produce these modulation parameters, i.e. $\gamma$ and $\beta$, from the vehicle’s real-time state dynamics. The encoder takes a time-series of sensor inputs, specifically the current state $\boldsymbol{x}_{t}$ and the previous state $\boldsymbol{x}_{t-1}$ as its input. Its operation begins by explicitly modeling the velocity-dependent force of damper. It estimates the discrete-time “velocity” of the state by calculating the difference between consecutive time steps as $\Delta\boldsymbol{x}_{t}=\boldsymbol{x}_{t}-\boldsymbol{x}_{t-1}$.
@@
$\gamma_{t},\beta_{t}]=\text{Linear}(g_{t}\odot D_{t}).$ (19)
III Experiments and Results
III-A Experimental Setup
@@
III-B Comparative Experiments
III-B1 Simulation Results
We select 10 simulation scenarios, including two different driving styles: smooth driving and aggressive driving, and test with 14 vehicle types. For each vehicle type, several 20-second data segments are collected under each scenario. Training set accounts for 70%. Sensor data required by the algorithm are directly obtained from the software, and the ground truth of wheel loads is also directly obtained. We choose LSTM-GA [1], E-PINN [6] and PINN [10] as comparative baseline methods, with RMSE (root mean square error) and MaxError (mean of max error of each segment) as evaluation metrics. The quantitative results are shown in Table I.
@@ III-B2 Real-World Test Results
III-C Ablation Studies
We conduct ablation experiments on several key modules, and the results are shown in Table III. It is noted here that “Basic model” refers to establishing a model for the entire vehicle [25] instead of modeling the linkage structure of the suspension.
@@ III-C Ablation Studies
Based on the results of the ablation experiments, when a Basic Model without suspension modeling is selected, due to the highly nonlinearity of the vehicle system, relying solely on the overall parameters of the vehicle body will lead to significant deviations. In addition, the guidance of the physical loss ensures that the output of the neural network meets the basic constraints and will not produce an excessively large MaxError. The denoising function of the Bayesian operation has also been proven to be effective, while NS-Dropout further improves the network performance.
IV Conclusion
This paper proposes a framework for measuring the dynamic wheel load of vehicles. Firstly, through refined modeling of the suspension linkage system, a nonlinear instantaneous dynamic system is established to characterize the complex geometric structure of the suspension. Subsequently, the chassis dynamics is integrated into the neural network as physical information by the DPC module, and Bayesian operations are adopted to enhance the network’s robustness against noisy inputs. We conduct experiments using CarSim simulation software and a Formula Student race car, verifying the effectiveness and stability of our method under various working conditions. However, considering the requirements for race car light-weighting and the constraints on on-board computing power, we have not yet realized real-time on-board computation. In addition, due to the need for special design of relevant chassis sensors, we have only conducted tests on the Formula Student race car for the time being and have not verified the method on passenger vehicles. Future work will attempt to solve the above issues.

### v3

@@ Abstract
Accurate state estimation is fundamental to intelligent vehicles. Wheel load, one of the most important chassis states, serves as an essential input for advanced driver assistance systems (ADAS) and exerts a direct influence on vehicle stability and safety. However, wheel load estimation remains challenging due to the complexity of chassis modeling and the susceptibility of nonlinear systems to noise. To address these issues, this paper first introduces a refined suspension linkage-level modeling approach that constructs a nonlinear instantaneous dynamic model by explicitly considering the complex geometric structure of the suspension. Building upon this, we propose a damper characteristics-based Bayesian physics-informed neural network (Damper-B-PINN) framework to estimate dynamic wheel load, which leverages the suspension dynamics as physical guidance of PINN while employing Bayesian inference to mitigate the effects of system noise and uncertainty. Moreover, a damper-characteristic physics conditioning (DPC) module is designed for embedding physical prior. The proposed Damper-B-PINN is evaluated using both high-fidelity simulation datasets generated by CarSim software and real-world datasets collected from a Formula Student race car. Experimental results demonstrate that our Damper-B-PINN consistently outperforms existing methods across various test conditions, including extreme ones. These findings highlight the potential of the proposed Damper-B-PINN framework to enhance the accuracy and robustness of dynamic wheel load estimation, thereby improving the reliability and safety of ADAS applications.
Figure 1: Real-world testing platform (A Formula Student racing car) and standard testing circuits.
@@ I Introduction
State estimation in real-world engineering systems subject to noise such as wheel load estimation in vehicle chassis systems, remains a longstanding challenge due to the complex and imperfect nonlinear relationships between inputs and outputs [36, 35, 29]. Traditional methods like EKF sometimes fail to achieve accurate estimation in such real-world scenarios [13]. In recent years, data-driven methods have increasingly been employed to approximate closures in nonlinear models and to estimate parameters and functional relationships within mathematical frameworks [30, 31]. Deep learning approaches have shown promise for handling such problems, with notable success in solving partial differential equations (PDEs) from sparse or noisy data [24]. Among these, physics-informed neural networks (PINNs) have emerged as a powerful paradigm for both forward (inference) and inverse (identification) problems due to their ability to integrate physical laws into the learning process [5, 28]. By embedding governing equations into the loss function, PINNs can infer unknown parameters and reconstruct solutions from partial observations, making them well-suited for complex state estimation tasks in nonlinear dynamical systems [14].
Recent studies on PINNs have explored their applications in various complex engineering fields, such as robotics [33], automotive [12], and transportation [15], demonstrating strong performance across diverse tasks. Additionally, novel optimization methods have been proposed to enhance the accuracy and generalization of PINNs [4]. Despite the fact that PINNs have demonstrated success in integrating physics into deep neural network (DNN) frameworks, enabling the concurrent use of physics as explicit knowledge and data as implicit knowledge, they lack built-in uncertainty quantification, limiting their applicability in scenarios with high noise levels.
Recently, physics-informed BNNs have been introduced as a hybrid approach that integrates PINNs and Bayesian techniques to quantify uncertainties in both data and models [6]. Considering sparse and noisy measurements in boundary conditions and source terms in real-world applications, Yang et al. [32] solved both forward and inverse problems involving linear and nonlinear PDEs with noisy data.
In the wheel load estimation task, there are two main challenges: (1) Inaccurate dynamic modeling will mislead the PINN, ultimately resulting in the physical model failing to match the data distribution. (2) The chassis sensors have significant noise, and traditional PINN is only constrained through the physical loss, making it fail to stably embed the physical information during the training process. To address these problems, we design a damper characteristics-based Bayesian physics-informed neural network (Damper-B-PINN), a BNN framework inspired by damper properties and introduce a physical model as a priori knowledge within the network structure to enhance accuracy in system estimation, as shown in Figure 2. The main contributions of this paper are summarized as follows:
  • A Bayesian PINN framework with a damper-characteristic physics conditioning (DPC) module, which is introduced to reduce the impact of environmental noise on systems with deeply embedded physics information, enhancing the robustness of networks.
@@
II Related Works
II-A Physics-Informed Neural Networks
The baseline PINN algorithms have achieved remarkable success in the solution of many linear and nonlinear PDEs. For instance, PINNs have been utilized for modeling cardiovascular flows, constraining the output to satisfy physical conservation laws through one-dimensional models of pulsatile blood flow [9]. The model proposed in [17] encoded the physical principles of the hydraulic fracturing process, described in the form of PDEs, in DNNs. Meng et al. [16] combined the first-order reliability method with PINNs to avoid calculating the real structure response, which tackled computational difficulty and inefficiency. Recently, many new PINN architectures, such as conservative PINNs [7] and nonlocal PINNs [20], have been proposed to improve the robustness and efficiency of PINNs. However, current PINN-based methods often lead to non-convergence issues and low accuracy due to system uncertainty and nonlinearity. Consequently, the application of PINNs in reliability assessment remains under-explored.
II-B Physics-Informed Bayesian Neural Networks
In the context of power systems, the robustness of physics-informed BNNs against noise-induced uncertainty for system identification has been evaluated, outperforming parsimonious approaches such as sparse identification of nonlinear dynamics [27]. PINNs have also been employed as a surrogate model to solve Bayesian inverse problems, significantly reducing computational time by retaining initial update results to ensure the global nature of the surrogate model [18]. In another work, an offline-online computational strategy that coupled classical sampling methods with PINN-based approaches for Bayesian inverse problems achieved a substantial reduction in overall computational time while maintaining accuracy [11]. Physics-informed BNNs and ensembled PINNs have also been leveraged to quantify uncertainties arising from noisy and incomplete data in governing equations, providing reasonable uncertainty bounds and improving the reliability of predictions [38]. Besides, Li et al. [10] designed a physics-informed deep learning framework for forward simulation and unknown parameter inference in nondiffusive thermal modeling using sparse and noisy data. However, existing physics-informed BNN-based models have not been thoroughly validated for complex real-world engineering problems, such as real-time dynamic wheel load estimation.
III Methodology
III-A Problem Formulation
@@
III-B Suspension Linkage-Level Modeling
Since vehicle suspensions typically consist of multiple connecting rods forming complex single-degree-of-freedom (1 DOF) or two-degree-of-freedom (2 DOF) structures, their geometric parameters evolve dynamically during driving [34]. For computational simplicity, the mass of individual suspension links is neglected, and only the unsprung mass $m_{u}$, which includes the wheel and steering knuckle, is considered. The hinge joints of the suspension system, are commonly referred to as “hard points”. Detailed components are labeled in Figure 3. In this study, the hard points are defined as follows: $\mathbb{u_{1}}$ and $\mathbb{u_{2}}$ (front and rear points of the upper control arm, UCA); $\mathbb{l_{1}}$ and $\mathbb{l_{2}}$ (front and rear points of the lower control arm, LCA); $\mathbb{p_{1}}$ and $\mathbb{p_{2}}$ (lower and upper joints of the spring-damper assembly); $\mathbb{t}$ (hinge joint between the steering knuckle and steering rack); and $\mathbb{s_{1}}$, $\mathbb{s_{2}}$, and $\mathbb{s_{3}}$ (upper, front, and lower points of the steering knuckle).
@@ III-B Suspension Linkage-Level Modeling
where $f$ represents all the solving processes from Equation (LABEL:eq2) to Equation (7). Then, the complex suspension dynamics will be integrated into the BNN as physical guidance.
III-C Bayesian Physics-Informed Neural Network
III-C1 Physics-Informed Bayesian Framework
@@
III-C2 Variational Inference with Physics-Informed Likelihood
@@
III-C3 Normal-Sigmoid-Dropout Method
@@
III-D Damper-Characteristic Physics Conditioning Embedding
@@
The DPC encoder is designed to produce these modulation parameters, i.e. $\gamma$ and $\beta$, from the vehicle’s real-time state dynamics. The encoder takes a time-series of sensor inputs, specifically the current state $\boldsymbol{x}_{t}$ and the previous state $\boldsymbol{x}_{t-1}$ as its input. Its operation begins by explicitly modeling the velocity-dependent force of the damper. It estimates the discrete-time “velocity” of the state by calculating the difference between consecutive time steps as $\Delta\boldsymbol{x}_{t}=\boldsymbol{x}_{t}-\boldsymbol{x}_{t-1}$.
@@
$\gamma_{t},\beta_{t}]=\text{Linear}(g_{t}\odot D_{t}).$ (19)
IV Experiments and Results
IV-A Experimental Setup
IV-B Comparative Experiments
IV-B1 Simulation Results
We select 10 simulation scenarios, including two different driving styles: smooth driving and aggressive driving, and test with 14 vehicle types. For each vehicle type, several 20-second data segments are collected under each scenario. Training set accounts for 70% of total. Sensor data required by the algorithm are directly obtained from the software, and the ground truth of wheel loads is also directly obtained. We choose LSTM-GA [36], E-PINN [5], and PINN [15] as comparative baseline methods, with RMSE (root mean square error) and MaxError (mean of max error of each segment) as evaluation metrics. The quantitative results are shown in Table I.
@@ IV-B1 Simulation Results
IV-B2 Real-World Test Results
@@
IV-C Ablation Studies
We conduct ablation experiments on several key modules, and the results are shown in Table III. It is noted here that “Basic model” refers to establishing a model for the entire vehicle [26] instead of modeling the suspension linkage.
@@ IV-C Ablation Studies
When a Basic Model without suspension modeling is selected, due to the highly nonlinearity of vehicle system, it will lead to significant deviations through relying solely on the overall parameters of vehicle body. In addition, the guidance of the physical loss ensures that the output of the neural network meets the basic constraints and will not produce an excessively large MaxError. The denoising function of the Bayesian operation has also been proven to be effective, while NS-Dropout further improves the network performance.
V Conclusion
This paper proposes a framework for measuring the dynamic wheel load of vehicles. Firstly, through refined modeling of the suspension linkage system, a nonlinear instantaneous dynamic system is established to characterize the complex geometric structure of the suspension. Subsequently, the chassis dynamics is integrated into the neural network as physical information by the DPC module, and Bayesian operations are adopted to enhance the network’s robustness against noisy inputs. We conduct experiments using CarSim simulation software and a Formula Student race car, verifying the effectiveness and stability of our method under various working conditions. However, considering the requirements for race car light-weighting and the constraints on on-board computing power, we have not yet realized real-time on-board computation. In addition, due to the need for special design of relevant chassis sensors, we have only conducted tests on the Formula Student race car and have not verified the method on passenger vehicles. Future work will attempt to solve the above issues.